### v1

# Multi-FAct: Assessing Multilingual LLMs’ Multi-Regional Knowledge using FActScore

Abstract
Large Language Models (LLMs) are prone to factuality hallucination, generating text that contradicts established knowledge. While extensive research has addressed this in English, little is known about multilingual LLMs. This paper systematically evaluates multilingual LLMs’ factual accuracy across languages and geographic regions. We introduce a novel pipeline for multilingual factuality evaluation, adapting FActScore (Min et al., 2023) for diverse languages. Our analysis across nine languages reveals that English consistently outperforms others in factual accuracy and quantity of generated facts. Furthermore, multilingual models demonstrate a bias towards factual information from Western continents. These findings highlight the need for improved multilingual factuality assessment and underscore geographical biases in LLMs’ fact generation [^1] Source code and data available at: https://github.com/sheikhshafayat/multi-fact.
1 Introduction
Large Language Models (LLMs) are susceptible to factuality hallucination, a phenomenon in which the generated text contradicts established world knowledge (Huang et al., 2023; Zhang et al., 2023). Despite extensive research focusing on hallucination and factuality of LLMs in free-form generation, these endeavors have predominantly concentrated on English (Huang et al., 2023; Min et al., 2023; Mishra et al., 2024). Consequently, there exists a notable gap in our comprehension of the factual accuracy of LLMs when producing content in various languages. As highlighted by Kang et al. (2024), current metrics for detecting hallucination are inadequate in multilingual settings. The extent to which LLMs exhibit factual hallucination remains unclear across different languages. Similar to other capabilities, there may be a discernible decline in performance when LLMs are tasked with non-English contexts (Ahuja et al., 2023; Bang et al., 2023).
(a) Factuality Heatmap of GPT-3.5 Biography Generation in English
(b) Factuality Heatmap of GPT-3.5 Biography Generation in Korean
Figure 1: While GPT-3.5 in English knows about the presidents of all countries almost uniformly, its knowledge in Korean is much smaller and mostly limited to East Asia and North America. Here color represents FActScore (bluer is better).
In this paper, we address this gap by systematically evaluating the factual accuracy of multilingual LLMs across different languages and geographic regions. We explore the following research questions: R1: Do multilingual models, such as ChatGPT, exhibit uniform factual accuracy across all languages in free-form generation? R2: Do these models demonstrate higher factual precision for content that aligns closely with the language (e.g., discussing a president of a country whose official language is used for generation)?
To address these questions, we introduce a novel pipeline tailored for evaluating factuality in a multilingual setting. Our approach first adapts the FActScore (Min et al., 2023) to accommodate multiple languages, and we make this pipeline openly accessible as open-source. Using this pipeline, we then assess the factual accuracy of nine languages in biography generation tasks, ensuring geographical diversity through a curated list of subjects.
Through our analysis, we observe two significant findings. First, English consistently maintains an advantage in both factual accuracy and the quantity of generated facts compared to other languages when generating identical content. Second, content produced by multilingual language models tends to exhibit a stronger performance for factual information originating from Western regions, such as America and Europe, across the languages. These results emphasize the need for enhanced assessment methods in evaluating multilingual factual accuracy and underscore the geographical biases in the fact generation capabilities of LLMs.
2 Methodology
2.1 Task Selection
Min et al. (2023) select biography generation as their task because biographies contain factual information readily verifiable from existing sources like Wikipedia, and confirm through human evaluation that English Wikipedia is suitable for biography verification tasks.
Topic Selection
We select “National Leaders of Each Countr” as the topic for factuality measurement, as it is fairly common across various cultures and languages to have well-known and well-described national leaders such as presidents and prime ministers. Specifically, we choose to focus on the president or head of state in the year 2015. This year is randomly selected from those prior to the emergence of LLMs. To determine our samples of the topic, we select countries by compiling a list of the 20 most populous nations from each continent. [^2] America, Asia, Europe, and Africa; America consisted of North and South America; Australia and New Zealand were also grouped with America. The lists of countries and their corresponding leaders used in our experiments are detailed in Table 2 of the Appendix. We choose a total of 9 languages, divided among high, middle and low resource languages: English, German, French, Spanish, Arabic, Swahili, Chinese, Korean and Bengali. We make sure to include at least three languages that are widely spoken in each continent. [^3] Africa: Arabic, Swahili, and French; Europe: Spanish, French, German; Asia: Korean, Bengali, Chinese; America: Spanish and French. English is widely spoken all over the world, and kept as the baseline.
Figure 2: Pipeline of our methodology. For details of step three, fact verification, please refer to the original FActScore Min et al. (2023).
Figure 2: Pipeline of our methodology. For details of step three, fact verification, please refer to the original FActScore Min et al. (2023).
2.2 Multi-FAct Pipeline
We introduce a novel pipeline for automatically measuring factuality of a multilingual LLM for multi-regional knowledge. This pipeline evaluates the biographies of presidents or heads of 80 countries across 9 languages.
The pipeline as shown in Figure 2 is structured into three main stages: 1) Obtaining multilingual generations (§ 2.3), 2) Translating these facts into English using GPT-3.5 (§ 2.4), 3) Measuring Factuality(§ 2.5).
2.3 Generating Facts
Following  Min et al. (2023), we prompt the model to generate a biography of designated topics.
Model
We use GPT-3.5 (gpt-3.5-turbo-0613) and GPT-4 (gpt-4-1106-preview) for the generation of facts. All experiments were carried out between January and February 2024. The model temperature was set at 1.0.
Prompt
We translate the following prompt into 9 languages: Write a biography of {name}. The prompts were translated by GPT-4 and verified by native speakers’ manual evaluation. We also transliterate each person’s name into corresponding languages and validated by native speakers.
2.4 Translating Generated Facts
To compare the biographies written in the eight non-English languages, we translate the generated content from the original languages to English using GPT-3.5, specifically, gpt-3.5-turbo-0125. In principle, any translation model that can generate coherent translation would work. We do not use commercially available machine translators like Google Translate because manual evaluation revealed that they can not maintain the consistent gender of the person throughout the whole text.
We translate the generation and verify facts in English, rather than doing fact verification in corresponding languages with respective Wikipedia articles, for several reasons. Firstly, Wikipedia differs in size and scope for each language Wikipedia contributors (2024), which makes it difficult to compare cross-lingual factuality if the knowledge base is different. Moreover, some important steps in the original FActScore pipeline work best in English (e.g., RAG, NPM are designed primarily for English). We, however, do some quantitative analysis on the effect of LLM-based translation on FActScore evaluation and find that translation by GPT-3.5 does not affect FActScore of texts significantly (see appendix A.2).
2.5 Measuring Factuality
To evaluate the accuracy of the generated model’s response M, we use the FActScore metric. This involves breaking down the translated generation into atomic facts—short sentences conveying a single piece of information, as defined in the FActScore paper. The accuracy of these atomic facts, denoted as A, is verified against the corresponding English Wikipedia article, serving as the reference knowledge source C.
[…]
We replace all proprietary models of the original FActScore pipeline with open-source models, ensuring cost-effectiveness while maintaining the quality of our system. We decompose LLM responses M into atomic facts using Mistral-7B Jiang et al. (2023). The verification step also uses Mistral-7B along with RAG and NPM Min et al. (2022) for the best performance. The details statistics can be found in Appendix A.2.
3 Results
Figure 3: FActScore for each language of GPT-3.5 and GPT-4 for each language.
Figure 4: Number of correct and hallucinated facts by GPT-3.5 and GPT-4 for each language. The number on top of each bar is the FActScore . Note that in some languages ChatGPT generates shorter responses, means even though they might be generating not very good responses (or refusing to respond), their FActScore is higher.
3.1 Does the Factuality Differ Across Languages?
Figure 3 displays the outcomes of measuring factuality across various languages using FActScore as a metric, uncovering significant variations among languages.
English, Spanish, French, German, and Swahili exhibit notably higher FActScore for both GPT-3.5 and GPT-4. Figure 5 shows the distribution of FActScore by language. Note that even in GPT4, some languages have long tail distribution in FActScore .
(a) GPT-4
(b) GPT-3.5
Figure 5: FActScore Distributions by language between GPT-4 and GPT-3.5
Figure 4 compares factuality across languages through the average number of correct and hallucinated facts, highlighting a significant gap in the quantity of facts generated between English and other languages. Remarkably, even with similar FActScore as shown in Figure 3, English surpasses other languages in generating a higher number of correct facts, thus delivering more useful responses.
These observations emphasize the influence of output length differences on the number of correct and hallucinated facts across languages, despite similar FActScore values. English and other high-resource languages (e.g., Chinese, Spanish) typically produce more facts due to longer outputs. In contrast, low-resource languages like Bengali yield fewer facts because of shorter responses, even when their FActScore are comparable to English (0.58 for Bengali, 0.61 for English). This discrepancy further widens the factuality gap between low-resource and high-resource languages. Hence, the denormalization of output length in multilingual contexts suggests that evaluating factuality requires considering both FActScore and the number of correct facts for a thorough assessment.
Language Africa America Asia Europe Mean STD
Spanish 0.624 0.641 0.627 0.640 0.633 0.087
German 0.622 0.642 0.616 0.628 0.627 0.094
Swahili 0.643 0.637 0.585 0.610 0.619 0.096
English 0.607 0.632 0.613 0.607 0.615 0.086
French 0.595 0.655 0.609 0.594 0.613 0.110
Bengali 0.579 0.574 0.585 0.589 0.582 0.149
Arabic 0.493 0.559 0.485 0.522 0.515 0.167
Korean 0.481 0.501 0.490 0.476 0.487 0.208
Chinese 0.453 0.502 0.479 0.514 0.487 0.230
Table 1: Overall FActScore by language with mean and standard deviation (STD). The continents with the highest and second-highest values for each language are emphasized in bold and underlined, respectively.
Figure 6: Continental distribution of top 20 countries that had the highest FActScore in each language. The right way to interpret this plot is, say for German, we take the twenty most accurate presidents’ biographies and then look into which continents they belong to.
3.2 Does the Factuality Geographically Differ Across Languages?
In this section, we explore whether the factuality of models varies with the geographic distribution of the language under study. Since GPT-4 performs better in low resource language as seen in figure 3, from here on, our analysis is conducted using only the outputs from GPT-4.
(a) English
(b) Chinese
(c) Korean
(d) Arabic
Figure 7: Fine-grained geographic distribution of languages with the highest standard deviation (Chinese, Korean, Arabic) and the least standard deviation (English) is given [^5] Please note that we discarded South Africa and Uzbekistan from this analysis because they corresponded to only elements in their respective geographic categories, while other categories represent averages of two or more countries. Details of which countries belong to which category can be found in appendix A.1. The top bar represents the average number of hallucinated facts, and the bottom one denotes correct facts. The number on top of each bar represents FActScore .
Table 1 presents the overall FActScore by language and continent. Interestingly, languages with higher mean values tend to have lower standard deviations, indicating that these languages maintain a relatively uniform level of factual accuracy regardless of the geographical area. Western languages such as English, Spanish, and German notably score high across continents, while Chinese, although a high-resource language, exhibits low factual precision. Additionally, among the nine languages evaluated, six demonstrate their highest performance in content related to the American continent, highlighting a prevalent American-centric bias in GPT’s knowledge across most languages.
To conduct an in-depth analysis, we counted which continent the top-K FActScore for each language pertains to, thereby mapping out the distribution across continents. Figure 6 presents the findings for when K equals 20, representing the top 25 percentile of the dataset.
Across analyzed languages, America and Europe are the primary focal points, underscoring their global influence. Even in languages like Korean and Chinese, distinct from these continents, the most accurate outputs primarily relate to American and European topics. For Arabic, while the highest FActScore is linked to Africa, likely reflecting its geographical roots, the second highest pertains to America. This trend highlights a Western-centric bias in the model’s factual content distribution across languages.
3.3 Geographical Biases in Factual Precision Across Subregions
In this section, we examine how languages with notable variations in factuality across continents exhibit distinct geographical biases. Further analysis involves breaking down the continents into sub-regions to compare the number of correct facts and FActScore among three languages: Chinese and Korean, which exhibit he highest standard deviation (SD), and English, which demonstrates the lowest SD among the four continents, as detailed in Table 1 of § 3.2.
As shown in Figure 5, Chinese stands out in Eastern Asia, achieving the highest number of correct facts among all regions analyzed, indicating a pronounced bias towards its main linguistic area. Korean, also categorized in Eastern Asia, shows comparable levels of factuality in Eastern Asia to that of regions associated with America and Europe, underscoring geographical biases in the model’s factual precision. English, while displaying a bias towards North America, its main region, exhibits relatively even factuality across other regions. This analysis underscores the geographical biases in factuality, with languages showing preferential accuracy towards regions where they are predominantly spoken (Figures for the other language can be found in Figure 10 in Appendix A.3.)
3.4 Is there a correlation between FActScore across languages?
Figure 8: FActScore correlation matrix computed over all the FActScore values per language.
Figure 8 shows the language-wise distribution of FActScore grouped by continents. English, Spanish, French and German have the most correlation with each other and these languages have very similar FActScore . One interesting observation is that, Swahili, even though has very similar FActScore with English, Spanish, French, and German, (table 1) does not correlate with those languages as much. This likely reflects the fact that these two languages have slightly different knowledge distributions, which we also observed in figure 6.
4 Qualitative Analysis
Model Generations with Low FActScores
Qualitative analysis on model generations, and generations with low FActScore reveals some interesting patterns. Firstly, LLMs provide much more informative outputs in English than in other language, which is also quantified in figure 4. For example, in the biography of Park Geun-hye, the former president of South Korea, GPT-4 generates the following sentence in English: “Park Geun-hye studied electrical engineering at Sogang University in Seoul and later pursued further studies at the University of Grenoble in France” while when asked in Korean, it only mentions “She studied in France (…)”. This highlights the need to develop a metric that measures beyond simple factuality of each atomic sentence, but the informativeness, or specificity of it. Some generation failures in Korean (the outliers in Figure 5) reveal that GPT-4 suggests there’s no famous person by that name and proceeds to generate a biography template. Interestingly, the English version of GPT-4 never refuses to generate a biography of the leaders in our list.
Comparison of GPT-3.5 & GPT-4
The qualitative difference between GPT-3.5 and GPT-4 is also striking. For example, while GPT-3.5 struggles to even generate fluent and coherent texts in Bengali, GPT-4 Bengali generations are mostly fluent and accurate, however the generations were short and lacking specific details. One interesting example regarding difference between GPT-3.5 and GPT-4 comes from Arabic. In GPT-3.5, when asked about Pakistani former president Mamnoon Hussain, GPT-3.5 hallucinates a biography non-existent "ambitious young man from Saudi Arabia", while GPT-4 rightfully identifies and writes the biography of former Pakistani president. GPT-4 also refused to answer more often than GPT-3.5; For example, in Chinese, it refuses to answer the biography of Sushir Koirala and claims there is no famous person of this name, however goes on to generate a fictional biography of an Indian person (with a disclaimer). GPT-4, too, does not always refrain from answering, when it doesn’t know. For example, when asked about Ollanta Humala in Chinese, it goes on to describe Ollantay Tambo, an archeological site in Peru.
5 Discussion & Limitations
In this study, we introduce a unique approach to examining the distribution of geo-culturally based knowledge using atomic factuality in long-form LLM generation. We observe that LLMs exhibit a Western-centric bias, which remains evident even in languages from non-Western regions. This discovery prompts intriguing inquiries regarding the representation of multilingual knowledge within LLMs. For instance, do LLMs learn to represent more facts about South America in Bengali by being exposed solely to content in English or Spanish? Or, to what extent is multilingual data required to acquire factual knowledge in a specific language? However, due to the proprietary nature of models like GPT-4, answering these questions proves to be challenging. Future research should delve deeper into this issue with language models that also have open datasets. Recent studies such as Dolma Soldaini et al. (2024) and OLMo Groeneveld et al. (2024) represent a step forward in this regard, although their focus is limited to monolingual English corpora and models.
One limitation of our current study pertains to small sample bias. We examine only a single sample from each country, which may not provide a comprehensive representation, especially considering that automated FActScore sometimes yields slight overestimations or underestimations. Consequently, caution is needed when interpreting individual scores. For instance, assertions such as "In Korean, the biography of the prime minister of Malaysia is 5% more accurate than that of the Romanian president" should be made with caution. In this paper, we present all findings as averages across languages, continents, or regions to mitigate sample bias. Future research should include multiple individuals, not only political leaders, from each country to alleviate this limitation.
Another limitation of our methodology is the varying durations national leaders have been in power, potentially biasing internet corpora in their favor. One solution could be to prompt models to generate a “country biography” or “capital biography” (e.g., “Write me an essay on Addis Ababa”). However, we opted not to pursue this in our study, as the original FActScore research did not include a human evaluation baseline for non-biographical domains.
6 Related Work
LLM Factuality Evaluation
The assessment of factual accuracy in natural language texts predates the advent of LLMs Guo et al. (2022). Our Multi-FAct pipeline follows the previous fact evaluation research, which involves validating claims based on external resource such as Wikipedia article Thorne et al. (2018); Krishna et al. (2022); Zhong et al. (2019) or Google search Chern et al. (2023). Our work is heavily inspired and based on Min et al. (2023), which suggests to measure the factuality of a long-form generated output of LLMs by breaking it down into atomic facts. Another line of work focuses on evaluating factual accuracy solely based on model output or internal states, eliminating the need for external knowledge bases like Wikipedia Manakul et al. (2023); Azaria and Mitchell (2023); Dhuliawala et al. (2023). While such methods offer resource efficiency, they often sacrifice accuracy compared to approaches leveraging external resources for fact verification.
Multilingual Factuality Evaluation
In the realm of multilingual factuality, Gupta and Srikumar (2021) presents the X-Fact dataset for multilingual fact-checking. Similarly, Thorne et al. (2018) have developed frameworks for multilingual fact-extraction and verification evaluation. Additionally, Aharoni et al. (2022) introduce mFACE, a dataset aimed at evaluating multilingual factual consistency. While these contributions significantly advance the understanding of multilingual factuality, they primarily focus on the assessment of factual accuracy in existing articles or benchmarks rather than examining the factual accuracy in the generative outputs of models. In this vein, Kang et al. (2024) explore the efficacy of existing metrics designed to detect hallucinations of LLM-generations, proposed for English texts, in a multilingual setting and point out they fall short in multilingual contexts. Our study extends the endeavor of developing a factuality evaluation metric that can be applied in multiple languages by introducing a Multi-Fact pipeline.
Geo-culture biases of LLMs
There has been considerable research regarding cultural and linguistic bias in LLMs Hovy and Yang (2021); Cao et al. (2023); Hershcovich et al. (2022); Huang and Yang (2023); Jin et al. (2023). A concurrent work Mirza et al. (2024) examines geographic knowledge gaps in LLMs and finds that GPT models show higher factuality in the Global North over the Global South, consistent with our findings. Manvi et al. (2024) also reach a similar conclusion. However, none of the studies concerns free-form generation with non-English languages.
7 Conclusion & Future Work
In this paper, we introduce a simple pipeline to evaluate factuality of LLMs in multilingual settings and demonstrate its effectiveness with the national leaders’ biographies in 80 countries around the world. Our findings reveal that multilingual models, such as GPT-4, tend to produce more factual content in high-resource languages, with notably longer outputs in English. Additionally, we observe a pervasive Western-centric bias across all languages; on average, even Asian and African languages demonstrate better performance for American and European leaders. However, a more detailed geographical analysis indicates that languages still reflect some aspects of their local geographic distribution, albeit to a lesser degree.
While factuality might be an interesting way to look into a model’s knowledge about a certain person, our current implementation treats every fact equally. FActScore does not differentiate between the informational value of facts. For instance, the statements “Person X attended MIT” and “Person X attended a famous university” are treated as equivalent, though the former is obviously more informative and useful. Future research should aim to distinguish between specific, valuable facts and generic, less informative ones.
Another area for future investigation could involve examining the consistency of model-generated facts across different languages; for example, whether the model’s depiction of Barack Obama’s childhood in English aligns with its portrayal in Chinese, or whether it keeps its focus solely on his political endeavors.
Appendix A Appendix
A.1 List of Countries and Presidents
Table 2 lists all the names of the countries, presidents and their Geographic locations. Please note that, we changed Australia and New Zealand to Oceania in the UN geoscheme for plotting purpose.
Country Continent Region President
Ethiopia Africa Eastern Africa Hailemariam Desalegn
Tanzania Africa Eastern Africa Jakaya Kikwete
Kenya Africa Eastern Africa Uhuru Kenyatta
Uganda Africa Eastern Africa Yoweri Museveni
Mozambique Africa Eastern Africa Filipe Nyusi
Madagascar Africa Eastern Africa Hery Rajaonarimampianina
DR Congo Africa Middle Africa Joseph Kabila
Angola Africa Middle Africa José Eduardo dos Santos
Cameroon Africa Middle Africa Paul Biya
Egypt Africa Northern Africa Abdel Fattah el-Sisi
Algeria Africa Northern Africa Abdelaziz Bouteflika
Sudan Africa Northern Africa Omar al-Bashir
Morocco Africa Northern Africa Abdelilah Benkirane
South Africa Africa Southern Africa Jacob Zuma
Nigeria Africa Western Africa Muhammadu Buhari
Ghana Africa Western Africa John Mahama
Côte d’Ivoire Africa Western Africa Alassane Ouattara
Niger Africa Western Africa Mahamadou Issoufou
Burkina Faso Africa Western Africa Michel Kafando
Mali Africa Western Africa Ibrahim Boubacar Keïta
Australia America Australia and New Zealand Tony Abbott
New Zealand America Australia and New Zealand John Key
Haiti America Caribbean Michel Martelly
Cuba America Caribbean Raúl Castro
Dominican Republic America Caribbean Danilo Medina
Mexico America Central America Enrique Peña Nieto
Guatemala America Central America Otto Pérez Molina
Honduras America Central America Juan Orlando Hernández
Nicaragua America Central America Daniel Ortega
United States America Northern America Barack Obama
Canada America Northern America Justin Trudeau
Brazil America South America Dilma Rousseff
Colombia America South America Juan Manuel Santos
Argentina America South America Cristina Fernández de Kirchner
Peru America South America Ollanta Humala
Venezuela America South America Nicolás Maduro
Chile America South America Michelle Bachelet
Ecuador America South America Rafael Correa
Bolivia America South America Evo Morales
Paraguay America South America Horacio Cartes
Uzbekistan Asia Central Asia Islam Karimov
China Asia Eastern Asia Xi Jinping
Japan Asia Eastern Asia Shinzo Abe
South Korea Asia Eastern Asia Park Geun-hye
Indonesia Asia South-Eastern Asia Joko Widodo
Philippines Asia South-Eastern Asia Benigno Aquino III
Vietnam Asia South-Eastern Asia Trng Tn Sang
Thailand Asia South-Eastern Asia Prayut Chan-o-cha
Myanmar Asia South-Eastern Asia Thein Sein
Malaysia Asia South-Eastern Asia Najib Razak
India Asia Southern Asia Narendra Modi
Pakistan Asia Southern Asia Mamnoon Hussain
Bangladesh Asia Southern Asia Sheikh Hasina
Iran Asia Southern Asia Hassan Rouhani
Afghanistan Asia Southern Asia Ashraf Ghani
Nepal Asia Southern Asia Sushil Koirala
Turkey Asia Western Asia Recep Tayyip Erdoğan
Iraq Asia Western Asia Fuad Masum
Saudi Arabia Asia Western Asia Salman of Saudi Arabia
Yemen Asia Western Asia Abdrabbuh Mansur Hadi
Russia Europe Eastern Europe Vladimir Putin
Ukraine Europe Eastern Europe Petro Poroshenko
Poland Europe Eastern Europe Andrzej Duda
Romania Europe Eastern Europe Klaus Iohannis
Czech Republic Europe Eastern Europe Miloš Zeman
Hungary Europe Eastern Europe János Áder
Belarus Europe Eastern Europe Alexander Lukashenko
Bulgaria Europe Eastern Europe Rosen Plevneliev
United Kingdom Europe Northern Europe David Cameron
Sweden Europe Northern Europe Stefan Löfven
Italy Europe Southern Europe Sergio Mattarella
Spain Europe Southern Europe Mariano Rajoy
Greece Europe Southern Europe Prokopis Pavlopoulos
Portugal Europe Southern Europe Marcelo Rebelo de Sousa
Germany Europe Western Europe Joachim Gauck
France Europe Western Europe François Hollande
Netherlands Europe Western Europe Mark Rutte
Belgium Europe Western Europe Charles Michel
Austria Europe Western Europe Heinz Fischer
Switzerland Europe Western Europe Simonetta Sommaruga
Table 2:
A.2 Multi-FAct Implementation Details
We replicated the original FActScore implementation and replaced InstructGPT for atomic fact generation with Mistral-7B. We also used Mistral 7B for the retrieval-based verification step. The final version uses retrieval + Mistral-7B + NPM. NPM threshold was set to 0.3.
A.2.1 Replication of original FActScore
We provide a comparison between our implementation with the results reported in the paper. For all the following experiments, we used a subset of ChatGPT-generation human-annotated factual data from the original paper.
Metric Human FS FS (ours) Error (ours) Error (Original)
Value 0.626 $\pm$ 0.238 0.637 $\pm$ 0.207 -0.011 $\pm$ 0.129 -0.0383 $\pm$ 0.121
Table 3: Summary of Replication. Error refers to the difference between human FActScore and FActScore determined by our method. Error (Original) refers to the error by the best-performing method reported on the paper. We also report the standard deviation alongside each metric.
Note that the Human FActScore value in the table 3 is different from what is provided in the paper as we averaged over our subset only. We replicated the best-performing method (retrieval + NPM) on this dataset reported in the paper and marked it as Error (Original). Our higher accuracy is likely due to Mistral 7B’s better capabilities and our better hyperparameter tuning.
Please note that there is some stochasticity in the automated evaluation of FActScore , which is why we need to be cautious of individual scores, and only work with scores averaged over a large number of samples.
A.2.2 Evaluation of Multi-FAct Pipeline
The basis of our pipeline relies crucially on the GPT3.5-based translation (Step 3 in figure 2). It is imperative to see whether this translation step hurts the estimation of FActScore or not.
Evaluation Method:
The basis of our evaluation method is that the factuality of a text is translation invariant, ie, factual knowledge in a text should not change when it is translated to other languages. To evaluate this, we first take the human-annotated examples provided in original FActScore , and translate them into four languages666one language from one continent using GPT4. We use GPT4 as previous studiesJiao et al. (2023) revealed GPT4 to be a good translator and better than GPT3.5 and commercially available machine translators. Native speakers of a subset of our languages also confirmed this.
Language/Dataset FActScore
Human-eval (EN) 0.626 $\pm$ 0.238
EN 0.637 $\pm$ 0.207
KOR 0.625 $\pm$ 0.211
BN 0.600 $\pm$ 0.216
EN 0.637 $\pm$ 0.207
FR 0.638 $\pm$ 0.200
SW 0.654 $\pm$ 0.196
Table 4: FActScore evaluation Across Languages. EN refers to FActScore evaluation on original English human annotation; other languages refer to FActScore evaluation on GPT4 translation of the human annotation.
We then applied our multilingual FActScore pipeline (Figure 2) on this GPT4 translated text. The premise is that if our pipeline (which involves GPT3.5 translation [^7] We did not use GPT4 translation in Multi-FAct pipeline to reduce costs) provides similar FActScore as the original English annotated data, it might provide good evidence that having a translation step in our pipeline doesn’t hurt FActScore estimation significantly.
Figure 9: Correlation of FActScore after on translated human evaluation data.
The results are provided in Table 4. We see that running Multi-FAct on non-English data does not significantly degrade FActScore evaluation.
We also note on the bottom row of Figure (3) that the correlation between FActScore evaluated in English and FActScore evaluated in other languages for the texts are very high; and all of them are very closely correlated with human-judgement.
A.3 Additional Results
Regional Breakdown (GPT4):
Figure 10 shows the regional breakdown for the other languages that were not included in Figure 5.
(a) Spanish
(b) German
(c) Swahili
(d) French
(e) Bengali
Figure 10: Fine-grained Geographic Distribution for Remaining Languages

### v2

@@
 A.2.2 Evaluation of Multi-FAct Pipeline
-The basis of our pipeline relies crucially on the GPT3.5-based translation (Step 3 in figure 2). It is imperative to see whether this translation step hurts the estimation of FActScore or not.
+The basis of our pipeline relies crucially on the GPT3.5-based translation (Step 3 in figure 2).
+It is imperative to see whether this translation step hurts the automated estimation of FActScore in a multilingual setting.
 Evaluation Method:
 The basis of our evaluation method is that the factuality of a text is translation invariant, ie, factual knowledge in a text should not change when it is translated to other languages.
-To evaluate this, we first take the human-annotated examples provided in original FActScore , and translate them into four languages666one language from one continent using GPT4.
+To evaluate this, we first take the human-annotated examples provided in original FActScore , and translate them into four languages using GPT4.
 We use GPT4 as previous studies Jiao et al.
 (2023) revealed GPT4 to be a good translator and better than GPT3.5 and commercially available machine translators.
-Native speakers of a subset of our languages also confirmed this.
+Manual inspection by native speakers of a subset of our languages also confirmed this.
 Language/Dataset FActScore
-Human-eval (EN) 0.626 $\pm$ 0.238
+Human eval (EN) 0.626 $\pm$ 0.238
 EN 0.637 $\pm$ 0.207
@@
 BN 0.600 $\pm$ 0.216
-EN 0.637 $\pm$ 0.207
 FR 0.638 $\pm$ 0.200
@@
 EN refers to FActScore evaluation on original English human annotation; other languages refer to FActScore evaluation on GPT4 translation of the human annotation.
+Figure 9: Correlation of FActScore evaluated on translated human-annotated data.
 We then applied our multilingual FActScore pipeline (Figure 2) on this GPT4 translated text.
-The premise is that if our pipeline (which involves GPT3.5 translation [^7] We did not use GPT4 translation in Multi-FAct pipeline to reduce costs) provides similar FActScore as the original English annotated data, it might provide good evidence that having a translation step in our pipeline doesn’t hurt FActScore estimation significantly.
-Figure 9: Correlation of FActScore after on translated human evaluation data.
-The results are provided in Table 4. We see that running Multi-FAct on non-English data does not significantly degrade FActScore evaluation.
+The premise is that if our pipeline (which involves GPT3.5 translation [^6] We did not use GPT4 translation in Multi-FAct pipeline to reduce API costs) provides similar FActScore as the original English annotated data, it might provide good evidence that having a translation step in our pipeline doesn’t hurt FActScore estimation significantly.
+The results are provided in Table 4. We see that running Multi-FAct on non-English data does not significantly degrade FActScore estimation.
 We also note on the bottom row of Figure (9) that the correlation between FActScore evaluated in English and FActScore evaluated in other languages for the texts are very high; and all of them are very closely correlated with human-judgement.

### v3

@@
-Multi-FAct: Assessing Multilingual LLMs’ Multi-Regional Knowledge using FActScore
+Multi-FAct: Assessing Factuality of Multilingual LLMs using FActScore
 Abstract
-Large Language Models (LLMs) are prone to factuality hallucination, generating text that contradicts established knowledge.
-While extensive research has addressed this in English, little is known about multilingual LLMs.
-This paper systematically evaluates multilingual LLMs’ factual accuracy across languages and geographic regions.
-We introduce a novel pipeline for multilingual factuality evaluation, adapting FActScore (Min et al., 2023) for diverse languages.
-Our analysis across nine languages reveals that English consistently outperforms others in factual accuracy and quantity of generated facts.
-Furthermore, multilingual models demonstrate a bias towards factual information from Western continents.
-These findings highlight the need for improved multilingual factuality assessment and underscore geographical biases in LLMs’ fact generation [^1] Source code and data available at: https://github.com/sheikhshafayat/multi-fact.
+Evaluating the factuality of long-form large language model (LLM)-generated text is an important challenge.
+Recently there has been a surge of interest in factuality evaluation for English, but little is known about the factuality evaluation of multilingual LLMs, specially when it comes to long-form generation.
+We introduce a simple pipeline for multilingual factuality evaluation, by applying FActScore (Min et al., 2023) for diverse languages.
+In addition to evaluating multilingual factual generation, we evaluate the factual accuracy of long-form text generation in topics that reflect regional diversity.
+We also examine the feasibility of running the FActScore pipeline using non-English Wikipediaand provide comprehensive guidelines on multilingual factual evaluation for regionally diverse topics.
 1 Introduction
 Large Language Models (LLMs) are susceptible to factuality hallucination, a phenomenon in which the generated text contradicts established world knowledge (Huang et al., 2023; Zhang et al., 2023).
-Despite extensive research focusing on hallucination and factuality of LLMs in free-form generation, these endeavors have predominantly concentrated on English (Huang et al., 2023; Min et al., 2023; Mishra et al., 2024).
-Consequently, there exists a notable gap in our comprehension of the factual accuracy of LLMs when producing content in various languages.
+Despite extensive research focusing on hallucination and factuality of LLMs in free-form generation, the previous works have predominantly studied English (Huang et al., 2023; Min et al., 2023; Mishra et al., 2024; Wei et al., 2024; Wang et al., 2023).
+Consequently, there exists a wide gap in our comprehension of the factual accuracy of LLMs when producing content in non-English languages.
 As highlighted by Kang et al.
@@
 Similar to other capabilities, there may be a discernible decline in performance when LLMs are tasked with non-English contexts (Ahuja et al., 2023; Bang et al., 2023).
-(a) Factuality Heatmap of GPT-3.5 Biography Generation in English
-(b) Factuality Heatmap of GPT-3.5 Biography Generation in Korean
-Figure 1: While GPT-3.5 in English knows about the presidents of all countries almost uniformly, its knowledge in Korean is much smaller and mostly limited to East Asia and North America.
-Here color represents FActScore (bluer is better).
 In this paper, we address this gap by systematically evaluating the factual accuracy of multilingual LLMs across different languages and geographic regions.
-We explore the following research questions: R1: Do multilingual models, such as ChatGPT, exhibit uniform factual accuracy across all languages in free-form generation?
-R2: Do these models demonstrate higher factual precision for content that aligns closely with the language (e.g., discussing a president of a country whose official language is used for generation)?
-To address these questions, we introduce a novel pipeline tailored for evaluating factuality in a multilingual setting.
-Our approach first adapts the FActScore (Min et al., 2023) to accommodate multiple languages, and we make this pipeline openly accessible as open-source.
-Using this pipeline, we then assess the factual accuracy of nine languages in biography generation tasks, ensuring geographical diversity through a curated list of subjects.
-Through our analysis, we observe two significant findings.
-First, English consistently maintains an advantage in both factual accuracy and the quantity of generated facts compared to other languages when generating identical content.
-Second, content produced by multilingual language models tends to exhibit a stronger performance for factual information originating from Western regions, such as America and Europe, across the languages.
-These results emphasize the need for enhanced assessment methods in evaluating multilingual factual accuracy and underscore the geographical biases in the fact generation capabilities of LLMs.
-2 Methodology
-2.1 Task Selection
-Min et al.
-(2023) select biography generation as their task because biographies contain factual information readily verifiable from existing sources like Wikipedia, and confirm through human evaluation that English Wikipedia is suitable for biography verification tasks.
-Topic Selection
-We select “National Leaders of Each Countr” as the topic for factuality measurement, as it is fairly common across various cultures and languages to have well-known and well-described national leaders such as presidents and prime ministers.
-Specifically, we choose to focus on the president or head of state in the year 2015.
-This year is randomly selected from those prior to the emergence of LLMs.
-To determine our samples of the topic, we select countries by compiling a list of the 20 most populous nations from each continent.
-[^2] America, Asia, Europe, and Africa; America consisted of North and South America; Australia and New Zealand were also grouped with America.
-The lists of countries and their corresponding leaders used in our experiments are detailed in Table 2 of the Appendix.
-We choose a total of 9 languages, divided among high, middle and low resource languages: English, German, French, Spanish, Arabic, Swahili, Chinese, Korean and Bengali.
-We make sure to include at least three languages that are widely spoken in each continent.
-[^3] Africa: Arabic, Swahili, and French; Europe: Spanish, French, German; Asia: Korean, Bengali, Chinese; America: Spanish and French.
-English is widely spoken all over the world, and kept as the baseline.
-Figure 2: Pipeline of our methodology. For details of step three, fact verification, please refer to the original FActScore Min et al. (2023).
-Figure 2: Pipeline of our methodology. For details of step three, fact verification, please refer to the original FActScore Min et al. (2023).
-2.2 Multi-FAct Pipeline
-We introduce a novel pipeline for automatically measuring factuality of a multilingual LLM for multi-regional knowledge.
-This pipeline evaluates the biographies of presidents or heads of 80 countries across 9 languages.
-The pipeline as shown in Figure 2 is structured into three main stages: 1) Obtaining multilingual generations (§ 2.3), 2) Translating these facts into English using GPT-3.5 (§ 2.4), 3) Measuring Factuality(§ 2.5).
+We explore the following research questions: R1: How can we effectively evaluate the factuality of multilingual LLMs in free-form text generation?
+R2: How do multilingual models’ free-form answers compare in factual accuracy across languages and geographically contextualized questions?
+To address these questions, we introduce Multi-FAct , a simple pipeline tailored for evaluating factuality in a multilingual context.
+Leveraging open-source models, we adapt the FActScore (Min et al., 2023) for multiple languages and validate it using human annotations from Min et al.
+(2023).
+Furthermore, we explore the use of non-English Wikipedia content to augment factuality evaluation, finding that while English Wikipedia remains a reliable source, concatenating articles from multiple non-English languages can sometimes provide estimates close to English Wikipedia.
+We conduct an evaluation of the factual accuracy of recent proprietary multilingual LLMs in nine languages using biography generation task.
+We ensure geographical diversity by curating subjects from various regions.
+Our analysis reveals two significant findings.
+First, English consistently maintains an advantage in both factual accuracy and the quantity of generated facts compared to other languages.
+Second, content produced by multilingual language models exhibits better factual accuracy about North America and Europe across the languages.
+Our contributions are as follows:
+  • We propose a novel pipeline Multi-FAct , that applies FActScore in a multilingual setting, and demonstrates the feasibility of conducting factuality evaluation of long-form generation in the multilingual biography task.
+  • We explore the potential and limitations of using non-English resources, particularly non-English Wikipedia articles, for factuality evaluation of long-form generation, showing that reasonably large non-English Wikipedia articles can also serve as a viable source.
+  • We introduce Multi-FAct as a tool to investigate cultural and geographic biases in LLMs, offering insights into how these biases manifest in LLM generated content.
+Figure 1: Our Multi-FAct Pipieline.
+The pipeline is structured into three main stages: 1) Obtaining multilingual generations, 2) Translating these facts into English using GPT-3.5, and 3) Measuring factuality, by breaking them down into smaller atomic facts and then verifying them by asking LLMs with context provided.
+2 Related Work
+LLM Factuality Evaluation
+The assessment of factual accuracy in natural language texts predates the advent of LLMs (Guo et al., 2022).
+Our Multi-FAct pipeline follows the previous fact evaluation research, which involves validating claims based on external resource such as Wikipedia article (Thorne et al., 2018; Krishna et al., 2022; Zhong et al., 2019) or Google search (Chern et al., 2023; Wei et al., 2024).
+Our work is inspired by and based on Min et al.
+(2023) which proposes FActScore to measure the factuality of a long-form generated output of LLMs by breaking it down into atomic facts.
+Another line of work focuses on evaluating factual accuracy solely based on model output or internal states, eliminating the need for external knowledge bases like Wikipedia (Manakul et al., 2023; Azaria & Mitchell, 2023; Dhuliawala et al., 2023).
+While such methods offer resource efficiency, they often sacrifice accuracy compared to approaches leveraging external resources for factuality evaluation.
+Multilingual Factuality Evaluation
+To evaluate factual accuracy of multilingual generation, Gupta & Srikumar (2021) presents the X-Fact dataset.
+Similarly, Thorne et al.
+(2018) presents frameworks for multilingual fact-extraction and verification evaluation.
+Additionally, Aharoni et al.
+(2022) introduces mFACE, a dataset aimed at evaluating multilingual factual consistency.
+While these papers advance the understanding of multilingual factuality, they focus on the assessment of factual accuracy in existing articles or benchmarks rather than examining the factual accuracy in the long-form generation of LLMs.
+Recently, Kang et al.
+(2024) explores the efficacy of existing metrics designed to detect hallucinations of LLM-generations in English in a multilingual setting, showing that the English-based metrics fall short in multilingual contexts.
+Our study extends the endeavor of developing a factuality evaluation metric that can be applied in multiple languages by introducing a Multi-Fact pipeline.
+Geo-culture biases of LLMs
+Research shows LLMs exhibit cultural and linguistic bias (Hovy & Yang, 2021; Cao et al., 2023; Hershcovich et al., 2022; Huang & Yang, 2023; Jin et al., 2023; Naous et al., 2023).
+A recent paper conducts a true-false evaluation task that shows GPT models are more accurate for facts about the Global North compared to the Global South (Mirza et al., 2024), consistent with our findings.
+We conduct a much more comprehensive study with long-form generation of geographically diverse information in multiple languages.
+3 Multi-FAct Pipeline
+We introduce Multi-FAct, a novel pipeline for automatically measuring factuality of a multilingual LLM in a multilingual setting.
+Multi-FAct evaluates the biographies of multi-regional topics in multiple languages.
+We choose the biography generation task because biographies are suitable for being decomposed into verifiable and independent "atomic" facts (Min et al.
+(2023)).
+The Multi-FAct pipeline is structured into three main steps: 1) Obtaining multilingual generations (§ 3.1), 2) Translating these facts into English using GPT-3.5 (§ 3.2), and 3) Measuring factuality of the translated facts (§ 3.3).
 3.1 Generating Facts
-Following  Min et al. (2023), we prompt the model to generate a biography of designated topics.
 Model
-We use GPT-3.5 (gpt-3.5-turbo-0613) and GPT-4 (gpt-4-1106-preview) for the generation of facts.
-All experiments were carried out between January and February 2024.
-The model temperature was set at 1.0.
+Most of the results in this paper use GPT4 (gpt-4-1106-preview) and GPT3.5 (gpt-3.5-turbo-0613) models because at the time of writing this paper, those were the only two strong multilingual models [^1] We also provide FActScore estimation for some newer models in Appendix C.
+for LLM multilingual biography generation.
+All experiments are conducted between January and March 2024222Except for GPT4o and GPT4o-mini in the Appendix., and the model temperatures are set to their defaults for respective models.
+Language
+We choose a total of nine languages: English, German, French, Spanish, Arabic, Swahili, Chinese, Korean, and Bengali.
+We carefully choose languages to represent the level of existing resources–high, medium, and low, and to represent diverse regions–Africa, Europe, Asia, and America  [^3] Africa: Arabic, Swahili, and French; Europe: Spanish, French, German; Asia: Korean, Bengali, Chinese; America: Spanish and French.
+English is widely spoken all over the world, and kept as the baseline..
 Prompt
-We translate the following prompt into 9 languages: Write a biography of {name}.
-The prompts were translated by GPT-4 and verified by native speakers’ manual evaluation.
-We also transliterate each person’s name into corresponding languages and validated by native speakers.
+We use the following prompt: Write a biography of {name}.
+We translate the prompt with GPT-4 into the eight non-English langauges, then we ask the native speakers to verify the translation.
+We also transliterate each person’s name into corresponding languages either using Wikipedia crosslinking or GPT-4, of which we take a subset and the native speakers verify that almost all the transliterations are correct.
 3.2 Translating Generated Facts
-To compare the biographies written in the eight non-English languages, we translate the generated content from the original languages to English using GPT-3.5, specifically, gpt-3.5-turbo-0125.
-In principle, any translation model that can generate coherent translation would work.
-We do not use commercially available machine translators like Google Translate because manual evaluation revealed that they can not maintain the consistent gender of the person throughout the whole text.
-We translate the generation and verify facts in English, rather than doing fact verification in corresponding languages with respective Wikipedia articles, for several reasons.
-Firstly, Wikipedia differs in size and scope for each language Wikipedia contributors (2024), which makes it difficult to compare cross-lingual factuality if the knowledge base is different.
-Moreover, some important steps in the original FActScore pipeline work best in English (e.g., RAG, NPM are designed primarily for English).
-We, however, do some quantitative analysis on the effect of LLM-based translation on FActScore evaluation and find that translation by GPT-3.5 does not affect FActScore of texts significantly (see appendix A.2).
+To compare the biographies written in the eight non-English languages, we translate the generated content from the original languages to English using gpt-3.5-turbo-0125.
+We do not use commercially available machine translation systems, such as Google Translate, based on the preliminary result that they do not maintain the consistent gender of the person throughout the text [^4] However, we do provide an analysis of what happens when Google Translation is used instead of GPT3.5 in Appendix C..
+We translate the generation and verify facts in English, rather than doing fact verification in corresponding languages with respective Wikipedia articles for multiple reasons.
+First, Wikipedia differs in size and scope for each language (Wikipedia contributors, 2024), which makes it difficult to compare cross-lingual factuality if the knowledge base is different.
+Additionally, key components of the original FActScore pipeline, such as RAG and NPM, are optimized for English and not available in other languages.
+Our quantitative analysis of LLM-based translation’s impact on FActScore evaluation reveals that GPT-3.5’s translation minimally influences the FActScore of texts (See § 3.4).
 3.3 Measuring Factuality
 To evaluate the accuracy of the generated model’s response M, we use the FActScore metric.
-This involves breaking down the translated generation into atomic facts—short sentences conveying a single piece of information, as defined in the FActScore paper.
-The accuracy of these atomic facts, denoted as A, is verified against the corresponding English Wikipedia article, serving as the reference knowledge source C.
+This involves breaking down the translated generation into atomic facts—short sentences, each conveying a single piece of information, as defined in Min et al.
+(2023).
+We denote the set of atomic facts as A, and we measure the accuracy of these atomic facts with the corresponding English Wikipedia article serving as the reference knowledge source C.
 […]
@@
 (2023).
-The verification step also uses Mistral-7B along with RAG and NPM Min et al.
-(2022) for the best performance.
-The details statistics can be found in Appendix A.2.
-3 Results
-Figure 3: FActScore for each language of GPT-3.5 and GPT-4 for each language.
-Figure 4: Number of correct and hallucinated facts by GPT-3.5 and GPT-4 for each language.
-The number on top of each bar is the FActScore .
-Note that in some languages ChatGPT generates shorter responses, means even though they might be generating not very good responses (or refusing to respond), their FActScore is higher.
-3.1 Does the Factuality Differ Across Languages?
-Figure 3 displays the outcomes of measuring factuality across various languages using FActScore as a metric, uncovering significant variations among languages.
+The verification step also uses Mistral-7B along with RAG and NPM for the best performance.
+We set the NPM threshold at 0.3.
+3.4 Reliability of Multi-FAct
+3.4.1 Replication of original FActScore
+We conduct a comparative analysis using our implementation, which includes a Mistral 7B model, instead of InstructGPT for atomic break-down, against the outcomes presented in the Min et al.
+(2023), using a subset of human-annotated factual data generated by ChatGPT (See Table 1).
+A comparison with other baselines and different parts of the Multi-FAct pipeline for different languages is provided in appendix A.
+Metric Human FS FS (ours) Error
+Value 0.626 $\pm$ 0.238 0.640 $\pm$ 0.127 -0.014 $\pm$ 0.128
+Table 1: Summary of Replication.
+Error refers to the difference between human FActScore and FActScore determined by our method.
+Standard deviation is reported alongside each metric.
+Note that the Human FActScore value is slightly different from what is provided in Min et al.
+(2023) as this value is averaged over our subset (44 samples).
+3.4.2 Effect of GPT-3.5 Translation in Multi-FAct
+As our pipeline crucially relies on the GPT-3.5 translation (Step 3 in figure 1), it is imperative to check whether the translation step affects the automated estimation of FActScore in the multilingual setting.
+We evaluate the effect of GPT-3.5 translation based on the premise that the factual content of a text should remain consistent across translations into different languages.
+In other words, if the factuality of an English text is established, its translated version in another language should exhibit the same level of factuality.
+Language Correlation Mean error Standard Deviation
+English 0.84 -0.014 0.13
+German 0.83 -0.007 0.13
+French 0.80 -0.004 0.14
+Spanish 0.81 -0.014 0.14
+Swahili 0.81 -0.015 0.13
+Arabic 0.82 0.021 0.14
+Chinese 0.81 0.040 0.14
+Korean 0.80 0.009 0.14
+Bengali 0.84 0.027 0.13
+Table 2: FActScore evaluation across languages.
+English refers to FActScore evaluation on original human annotation data provided in English; other languages refer to Multi-FAct evaluation on GPT-4 translation of the human annotation.
+Correlation refers to the correlation of individual FActScore with human-annotated FActScore provided in Min et al.
+(2023).
+To simulate and test the reliability of Multi-FAct with respect to translation, we first take the human-annotated examples provided in the original work (Min et al., 2023) and translate them into the eight non-English languages using GPT-4555We use GPT-4 as a previous study (Jiao et al., 2023) demonstrated GPT-4 to be a good translator..
+Then, we apply our multilingual FActScore pipeline (Figure 1) on the translated texts [^6] We did not use GPT-4 for back-translation in Multi-FAct pipeline to reduce API costs..
+Table 2 indicates that the execution of Multi-FAct on non-English data does not significantly degrade the estimation of FActScore, and the correlation between FActScore in different languages and human evaluation remains comparable.
+This outcome implies that adding a translation step to our pipeline does not significantly influence the estimation of FActScore .
+4 Reference Sources in Multilingual Factuality Evaluation
+(a) Number of total Wikipedia articles available per language (in millions).
+(b) Number of active Wikipedia users per language (note the log scale).
+(c) Number of original FActScore annotated topics that have Wikipedia articles in corresponding languages.
+Figure 2: The Wikipedia size distribution for languages in this study.
+(c) shows the number of human-annotated examples that also have representation in the corresponding Wikipedia.
+Note that Wikipedia size differs widely across languages, which usually means non-English Wikipedia articles are not as comprehensive as English Wikipedia articles for fact verification.
+We evaluate the factual accuracy of LLM-generated biographies against English Wikipedia for its reliability and comprehensive coverage of information about individuals, following prior works (Petroni et al., 2020; Chen et al., 2017; Min et al., 2023).
+However, we need to analyze the reliability of non-English Wikipedia articles as references.
+While the Multi-FAct pipeline is effective for topics with established English Wikipedia articles, its applicability may be limited because many regional topics (e.g., local politicians) lack English Wikipedia articles.
+If we cannot extend our Multi-FAct pipeline to these topics, it would be a major obstacle to ensuring LLM factuality with geographical diversity.
+However, as figure 2 illustrates, languages included in our study exhibit a wide range of variability in their representation within Wikipedia.
+In this section, we explore the following question: How should we measure multilingual factuality, when we cannot find a sufficient golden source of reference in English?
+First, we examine whether simple translation of Wikipedia articles from languages other than English would sustain the quality (§ 4.1).
+Then, we reveal that the "length of articles" is one factor contributing to performance degradation when non-English Wikipedia is used as fact verification ground truth corpus, and we propose a new method of concatenating articles from different languages to enhance the comprehensiveness of the reference (§ 4.2).
+4.1 Using Translated Wikipedia Articles as Reference
+We first test whether we can simply translate the Wikipedia articles written in eight non-English languages into English then replace the English Wikipedia articles within Multi-FAct pipeline with the translations.
+Using GPT-3.5, we translate the respective Wikipedia articles into English and conduct factual evaluation in English.
+We chose 157 individuals as topics for whom human-annotated data exists in Min et al.
+(2023), which served as the ground truth.
+We then compare this ground truth with the experimental results.
+Figure 3(a) illustrates the errors for each language’s Wikipedia against ground truth, and we include English Wikipedia corpus as the baseline.
+The results show a near-zero error rate for English, and lower error rates in French, Spanish, and German.
+We observe relatively higher error rates for Bengali, Korean, Chinese, and Arabic.
+Note that the size and coverage of Wikipedia varies across languages, thus the number of samples used in the experiments also varies by language, as indicated by $n$ in the figures.
+4.2 Effect of Reference Length
+(a) FActScore estimation error when using different Wikipedia corpus as ground truth.
+(b) FActScore estimation error scales linearly with the length of the Wikipedia article.
+Figure 3: Effect of using non-English Wikipedia as knowledge corpus for FActScore estimation.
+The x-axis in subfigure 3(b) represents the word count (in English) after translation.
+As the length of the non-English Wikipedia article increases, its utility for evaluating factuality increases.
+(a) Multi-FAct estimation for persons with ‘large’ Wikipedia articles in corresponding language.
+(b) Ensembling multiple non-English Wikipedias work better than single language Wikipedia.
+Figure 4: Effect of corpus size on FActScore Estimation. Figure 4(a) shows that comprehensive non-English sources might be used for factuality evaluation after translating into English.
+We hypothesize that non-English Wikipedias result in higher errors compared to English Wikipedia due to the lower coverage of topics, as evidenced by the shorter length of non-English articles.
+Figure 3(b) shows that there is a linear relationship between FActScore estimation error and the word count of Wikipedia articles, indicating that longer Wikipedia articles tend to serve better as references for factuality evaluation.
+To further test the linear relationship observed in Figure 3(b), we select 100 individuals from each of our studied languages who have both English Wikipedia articles and substantially long Wikipedia articles in their regional language.
+[^7] The detailed sampling method is provided in appendix C.2 We generate biographies for these 100 individuals in English and evaluate these biographies against both the English Wikipedia and the Wikipedia in the corresponding regional language.
+Since FActScore estimation for English is nearly perfect, we use it as a benchmark to estimate the error when using non-English articles as references.
+Figure 4(a) illustrates that the linear relationship between the length of Wikipedia articles and FActScore holds for longer articles as well.
+Ensembling Wikipedia
+Building upon the insights about the reference length discussed in § 4.1, this section investigates whether combining multiple non-English Wikipedias, referred to as “ensembling Wikipedias,” can enhance the accuracy of Multi-FAct measurements.
+In our experiments, we explore three cases: the union of wikis from Asian languages (Asian), the union of wikis from European languages (Euro), and the union of all wikis except English (All-Eng).
+As depicted in the Fig 3(a), among the two wikis with errors close to zero—namely, “Euro” and “All-Eng” cases—it is speculated that this could be attributed to their comprehensiveness or the high content overlap with the English Wikipedia.
+In the case of “Asian,” an error rate of 0.2 is observed, representing a decrease compared to using only one wiki of Asian language, and even approaching the performance observed for European languages (Spanish, French, and German).
+Our findings suggest that even for the same topic, the coverage of Wikipedia articles may vary across languages, thereby demonstrating that an “ensemble Wikipedia” comprising multiple articles can mitigate the performance gap between non-English Wikipedias and the English Wikipedia.
+However, the challenge of Wikipedia containing only a single, brief article remains unaddressed.
+5 Factual Accuracy of Multilingual models
+We use Multi-FAct to investigate the factual accuracy of multilingual LLMs in biography generation.
+We select “National Leaders of Each Country” as the topic for factuality measurement, as it is fairly common across various cultures and languages to have well-known and well-described national leaders such as presidents and prime ministers.
+Specifically, we choose to focus on the president or head of state in the year 2015.
+This year is randomly selected from those prior to the emergence of LLMs and also to make sure that at least some amount of data about the leader is online, which might not be true for very recent leaders.
+The lists of countries and their corresponding leaders used in our experiments are detailed in Table 8 of Appendix.
+5.1 Does factuality differ across languages?
+Figure 5: FActScore of GPT-3.5 and GPT-4 for each language. Note that GPT-4 makes a strong jump from GPT-3.5, especially in low-resource languages.
+Figure 6: Number of correct and hallucinated facts by GPT-3.5 and GPT-4 for each language. The number on top of each bar is the FActScore .
+Figure 6 illustrates the factuality across various languages using FActScore as a metric, uncovering significant variations among languages.
 English, Spanish, French, German, and Swahili exhibit notably higher FActScore for both GPT-3.5 and GPT-4.
-Figure 5 shows the distribution of FActScore by language.
-Note that even in GPT4, some languages have long tail distribution in FActScore .
-(a) GPT-4
-(b) GPT-3.5
-Figure 5: FActScore Distributions by language between GPT-4 and GPT-3.5
 Figure 6 compares factuality across languages through the average number of correct and hallucinated facts, highlighting a significant gap in the quantity of facts generated between English and other languages.
-Remarkably, even with similar FActScore as shown in Figure 3, English surpasses other languages in generating a higher number of correct facts, thus delivering more useful responses.
+Notably, even with similar FActScore as shown in Figure 6, English surpasses other languages in generating a higher number of correct facts, thus delivering more useful responses.
 These observations emphasize the influence of output length differences on the number of correct and hallucinated facts across languages, despite similar FActScore values.
@@
 Hence, the denormalization of output length in multilingual contexts suggests that evaluating factuality requires considering both FActScore and the number of correct facts for a thorough assessment.
+5.2 Does the Factuality Geographically Differ Across Languages?
+In this section, we explore whether the factuality of models varies with the geographic distribution of the language.
+Since GPT-4 performs better in low resource language (Figure 6), from here on, we conduct our analysis using only the outputs from GPT-4.
+Table 3 presents the overall FActScore by language and continent.
+Interestingly, languages with higher mean values tend to have lower standard deviations, indicating that these languages maintain a relatively uniform level of factual accuracy regardless of the geographical area.
+Western languages such as English, Spanish, and German notably score high across continents, while Chinese, although a high-resource language, exhibits low factual precision.
+Additionally, among the nine languages evaluated, six demonstrate their highest performance in content related to the American continent, highlighting a prevalent American-centric bias in GPT’s knowledge across most languages.
+Next, we analyze the geographic distribution of the most factually accurate information across different languages.
+For each language, we identify the continents associated with top-20 FActScore .
+Our findings reveal a strong bias towards American and European topics across all languages studied.
+This pattern persists even in languages like geographicallly distant from these continents, such as Korean and Chinese.
+For Arabic, while the highest FActScore is linked to Africa, likely reflecting its geographical roots, American topics still rank second place.
+This trend highlights a Western-centric bias in the model’s factual content distribution across languages.
+The result figure and detailed analysis of geographical biases across more fine-grained subregions are presented in Appendix D.
 Language Africa America Asia Europe Mean STD
@@
 Chinese 0.453 0.502 0.479 0.514 0.487 0.230
-Table 1: Overall FActScore by language with mean and standard deviation (STD).
+Table 3: Overall FActScore breakdown of GPT4 by language with mean and standard deviation (STD).
 The continents with the highest and second-highest values for each language are emphasized in bold and underlined, respectively.
-Figure 6: Continental distribution of top 20 countries that had the highest FActScore in each language.
-The right way to interpret this plot is, say for German, we take the twenty most accurate presidents’ biographies and then look into which continents they belong to.
-3.2 Does the Factuality Geographically Differ Across Languages?
-In this section, we explore whether the factuality of models varies with the geographic distribution of the language under study.
-Since GPT-4 performs better in low resource language as seen in figure 3, from here on, our analysis is conducted using only the outputs from GPT-4.
+Note that for all language except for Swahili, GPT4 is leaning towards American and European leaders in terms of factual accuracy.
+6 Conclusion
+In this paper, we introduce Multi-FAct , an extension of FActScore in multilingual setting to address the critical gap in factuality evaluation for non-English free-form generation from LLMs.
+We empirically demonstrate the viability of conducting factuality evaluation in the multilingual biography generation task.
+Our approach of translating generated content and comparing it against English Wikipedia proved effective, offering a scalable method for assessing factual precision across multiple languages.
+Additionally, we find that sufficiently comprehensive non-English sources can serve as alternatives to English sources using the same pipeline.
+Our analysis of LLMs’ outputs across different languages and geographically contextualized questions reveals a consistent advantage for English in terms of both factual accuracy and the quantity of generated facts, as well as better performance for content related to North America and Europe across languages.
+Multi-FAct serves as a valuable tool for investigating these biases, emphasizing the need for efforts to improve the cultural and geographic fairness of factual generations from LLMs.
+7 Limitations
+Our work shares some of the same limitations that original FActScore faces.
+Throughout the whole paper, we assume English Wikipedia as the most authoritative knowledge base for fact verification.
+However, it is known that Wikipedia itself has a Western-centric bias Naous et al.
+(2023).
+Moreover, when combining multiple non-English Wikipedia, we do not take into account when Wikipedia articles in different languages contradict each other, which can happen for controversial topics.
+It is also worth noting that FActScore and Multi-FAct both do not differentiate between the informational value of facts.
+For example, statements like “Person X attended MIT” and “Person X attended a renowned university” are treated equally, despite the former being more specific and informative.
+Future research should aim to distinguish between specific, valuable facts and generic, less informative ones.
+Acknowledgments
+This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.
+RS-2024-00509258, AI Guardians: Development of Robust, Controllable, and Unbiased Trustworthy AI Technology).
+We thank Pang Wei Koh and Sewon Min for providing valuable feedback on an earlier draft of this work.
+Appendix A Comparison with Baseline
+We provide a comparison of different components of Multi-FAct below.
+Trivial baselines: Random prediction: 8.86% Always "Supported": - 41.15% Always "Not Supported": 58.86%
+Method Arabic Swahili Bengali Chinese Korean French German Spanish Mean
+No Retrieval ChatGPT -0.268 -0.282 -0.260 -0.265 -0.277 -0.278 -0.288 -0.290 -0.276
+Retrieval + ChatGPT -0.163 -0.190 -0.165 -0.168 -0.177 -0.177 -0.175 -0.187 -0.175
+Retrieval + Mistral -0.150 -0.180 -0.169 -0.162 -0.172 -0.171 -0.175 -0.174 -0.169
+NPM -0.097 -0.133 -0.105 -0.096 -0.098 -0.140 -0.149 -0.153 -0.121
+Retrieval + Mistral + NPM 0.028 -0.016 0.022 0.029 0.006 -0.013 -0.013 -0.012 0.004
+Table 4: Multi-FAct pipeline component comparison across different languages.
+In Table 4:
+  • ChatGPT refers to GPT3.5-turbo.
+  • No Retrieval ChatGPT means simply asking ChatGPT whether a fact is correct or not (closed book setting).
+  • Retrieval + ChatGPT implies retrieving relevant passages from Wikipedia and then answering whether a fact is correct or not using GPT3.5.
+  • Retrieval + Mistral refers to retrieving passages and then using Mistral 7B to answer whether a fact is correct or not.
+This is similar to the previous setting except for the fact verification LLM is being switched to an open-source model.
+  • NPM refers to Non-Parametric Model (Min et al. (2022)), which requires having access to the Wikipedia article (ie, retrieval enabled).
+  • The final method is the ensemble version of NPM and Retrieval + Mistral and it performs the best (also reported in original FActScore work (Min et al.
+(2023))).
+This is the one we used throughout our paper for other analyses.
+While GPT3.5 and Mistral 7B are very similar at atomic fact breakdown, we use Mistral 7B in all experiments to save cost.
+Appendix B Effect of non-LLM Based Translation
+In section 3.4.2, we showed that GPT3.5-based translation does not hurt the FActScore estimate.
+We have also explained the reason we use GPT3.5 over commercial translation systems like Google Translate is because Google Translate is not able to keep the gender of the subject consistent across the generation.
+Table 5 provides a comparison of translation of FActScore estimation when the translation is done by Google Translate vs GPT3.5.
+Source Language GPT3.5 Error GTranslate Error
+Arabic 0.028 0.051
+Swahili -0.0157 0.0168
+Bengali 0.0219 0.057
+Chinese 0.0289 0.0883
+Korean 0.0063 0.0453
+French -0.0126 -0.0106
+German -0.0127 -0.0019
+Spanish -0.0115 0.0581
+Mean 0.0041 0.0380
+Table 5: FActScore estimation error rate comparison when translation to English was done using GPT-3.5 and GTranslate across different languages
+The experiment set up for Table 5 was the same as Table 2. The small difference in Error with GPT3.5 column arises from the run-to-run stochasticity of the FActScore estimation.
+Can we skip the translation step altogether?
+We also experimented with generating atomic facts in the target language (thus, skipping the translation step) with French and Korean using GPT3.5 Turbo.
+Qualitative analysis with native Korean speakers revealed that the generated atomic facts are quite reasonably good.
+However, because there is no NPM model available for Korean and French, the FActScore estimates using atomic facts generated with those methods do not match the best-performing method (Retrieval+Mistral+NPM) provided in the above table.
+Moreover, tokenizer fertility for non-English languages is often much worse than for English, which makes it much more computationally expensive to do atomic fact breakdown in Korean.
+At the time of writing this paper, no small open-source model, like Mistral 7B, could perform atomic fact breakdown in a diverse multilingual setting.
+Appendix C Additional Results
+C.1 Results from other models
+Table 6 shows the benchmark results from Claude-3 and Gemini-1.0-pro model. We chose to use GPT4 for most of our analysis because GPT4 has the best performance in low resource languages.
+Language (Claude) haiku opus sonnet gemini-1.0-pro gpt3.5 gpt4-turbo gpt4o-mini gpt4o
+English 0.68 0.70 0.62 0.44 0.57 0.62 0.54 0.58
+French 0.65 0.64 0.60 0.42 0.59 0.61 0.56 0.58
+Spanish 0.66 0.70 0.60 0.47 0.57 0.63 0.59 0.59
+German 0.63 0.66 0.58 0.43 0.57 0.63 0.55 0.60
+Chinese 0.27 0.43 0.19 0.40 0.40 0.49 0.49 0.53
+Korean 0.39 0.45 0.27 0.40 0.18 0.48 0.49 0.52
+Arabic 0.33 0.45 0.26 0.32 0.24 0.53 0.46 0.51
+Bengali 0.26 0.36 0.18 0.40 0.12 0.58 0.48 0.58
+Swahili 0.57 0.62 0.54 0.46 0.59 0.62 0.65 0.65
+Mean 0.49 0.56 0.43 0.42 0.43 0.58 0.53 0.57
+Table 6: FActScore estimation of the other models on the same national leaders’ biography generation task.
+Note that only GPT4-turbo has the overall highest performance and on average GPT4 series models have good performance across all languages.
+Note that model response size varies significantly across languages and models, which is not captured in this table.
+Similar to what is shown in figure 6, most of the models in this table tend to generate much longer responses for English and high-resource languages.
+The average response length also differs from model family to model family.
+C.2 Effect of Reference Length
+Sampling Method
+We randomly collected 10,000 people from 9 different Wikipedia using Wikipedia API and sorted them by their byte length.
+After that, we kept the top 100 entities that also had an English Wikipedia.
+Note that since we did this for each language separately, the list of 100 people differs for each language.
+However, manual inspection revealed that these entities are usually famous politicians, religious figures, famous athletes, and singers.
+Table 7 shows the word count distribution for the selected Wikipedia articles.
+Language Mean Text Length (Language Wiki) Mean Text Length (English Wiki)
+Spanish 3353.94 6895.68
+French 3617.13 6569.71
+German 3519.03 5723.78
+Korean 1520.79 5144.43
+Chinese 2879.44 4937.70
+Arabic 2829.48 5648.35
+Swahili 865.85 5003.25
+Bengali 2772.49 5829.74
+Table 7: Comparison of mean text lengths (in words) in Various Language Wikipedia after translating to English vs length of the same articles English Wikipedia.
+For the same subjects, English Wikipedia has much larger articles than the other languages.
+Appendix D Geographical Biases in Factual Precision Across Subregions
+Figure 7: Continental distribution of top 20 countries that had the highest FActScore in each language for GPT4.
+The way to interpret this plot is, say for German, we sort the twenty most accurate presidents’ biographies and then look up which continents they belong to.
+Note that for almost all languages, the top twenty biographies belong to subjects from America and Europe.
+We examine how languages with notable variations in factuality across continents exhibit distinct geographical biases.
+Further analysis involves breaking down the continents into sub-regions to compare the number of correct facts and FActScore among three languages: Chinese and Korean, which exhibit he highest standard deviation (SD), and English, which demonstrates the lowest SD among the four continents, as detailed in Table 3 of § 5.2.
+Chinese stands out in Eastern Asia, achieving the highest number of correct facts among all regions analyzed, indicating a pronounced bias towards its main linguistic area.
+Korean, also categorized in Eastern Asia, shows comparable levels of factuality in Eastern Asia to that of regions associated with America and Europe, underscoring geographical biases in the model’s factual precision.
+English, while displaying a bias towards North America, its main region, exhibits relatively even factuality across other regions.
+Figure 8 shows the fine-grained geographic distribution of languages with the highest standard deviation (Chinese, Korean, Arabic) and the least standard deviation (English).
 (a) English
@@
 (d) Arabic
-Figure 7: Fine-grained geographic distribution of languages with the highest standard deviation (Chinese, Korean, Arabic) and the least standard deviation (English) is given [^5] Please note that we discarded South Africa and Uzbekistan from this analysis because they corresponded to only elements in their respective geographic categories, while other categories represent averages of two or more countries.
-Details of which countries belong to which category can be found in appendix A.1.
+Figure 8: Fine-grained geographic distribution of languages with the highest standard deviation (Chinese, Korean, Arabic) and the least standard deviation (English) is given.
 The top bar represents the average number of hallucinated facts, and the bottom one denotes correct facts.
 The number on top of each bar represents FActScore .
-Table 1 presents the overall FActScore by language and continent.
-Interestingly, languages with higher mean values tend to have lower standard deviations, indicating that these languages maintain a relatively uniform level of factual accuracy regardless of the geographical area.
-Western languages such as English, Spanish, and German notably score high across continents, while Chinese, although a high-resource language, exhibits low factual precision.
-Additionally, among the nine languages evaluated, six demonstrate their highest performance in content related to the American continent, highlighting a prevalent American-centric bias in GPT’s knowledge across most languages.
-To conduct an in-depth analysis, we counted which continent the top-K FActScore for each language pertains to, thereby mapping out the distribution across continents.
-Figure 6 presents the findings for when K equals 20, representing the top 25 percentile of the dataset.
-Across analyzed languages, America and Europe are the primary focal points, underscoring their global influence.
-Even in languages like Korean and Chinese, distinct from these continents, the most accurate outputs primarily relate to American and European topics.
-For Arabic, while the highest FActScore is linked to Africa, likely reflecting its geographical roots, the second highest pertains to America.
-This trend highlights a Western-centric bias in the model’s factual content distribution across languages.
-3.3 Geographical Biases in Factual Precision Across Subregions
-In this section, we examine how languages with notable variations in factuality across continents exhibit distinct geographical biases.
-Further analysis involves breaking down the continents into sub-regions to compare the number of correct facts and FActScore among three languages: Chinese and Korean, which exhibit he highest standard deviation (SD), and English, which demonstrates the lowest SD among the four continents, as detailed in Table 1 of § 3.2.
-As shown in Figure 5, Chinese stands out in Eastern Asia, achieving the highest number of correct facts among all regions analyzed, indicating a pronounced bias towards its main linguistic area.
-Korean, also categorized in Eastern Asia, shows comparable levels of factuality in Eastern Asia to that of regions associated with America and Europe, underscoring geographical biases in the model’s factual precision.
-English, while displaying a bias towards North America, its main region, exhibits relatively even factuality across other regions.
-This analysis underscores the geographical biases in factuality, with languages showing preferential accuracy towards regions where they are predominantly spoken (Figures for the other language can be found in Figure 10 in Appendix A.3.)
-3.4 Is there a correlation between FActScore across languages?
-Figure 8: FActScore correlation matrix computed over all the FActScore values per language.
-Figure 8 shows the language-wise distribution of FActScore grouped by continents.
-English, Spanish, French and German have the most correlation with each other and these languages have very similar FActScore .
-One interesting observation is that, Swahili, even though has very similar FActScore with English, Spanish, French, and German, (table 1) does not correlate with those languages as much.
-This likely reflects the fact that these two languages have slightly different knowledge distributions, which we also observed in figure 6.
-4 Qualitative Analysis
-Model Generations with Low FActScores
-Qualitative analysis on model generations, and generations with low FActScore reveals some interesting patterns.
-Firstly, LLMs provide much more informative outputs in English than in other language, which is also quantified in figure 4.
-For example, in the biography of Park Geun-hye, the former president of South Korea, GPT-4 generates the following sentence in English: “Park Geun-hye studied electrical engineering at Sogang University in Seoul and later pursued further studies at the University of Grenoble in France” while when asked in Korean, it only mentions “She studied in France (…)”.
-This highlights the need to develop a metric that measures beyond simple factuality of each atomic sentence, but the informativeness, or specificity of it.
-Some generation failures in Korean (the outliers in Figure 5) reveal that GPT-4 suggests there’s no famous person by that name and proceeds to generate a biography template.
-Interestingly, the English version of GPT-4 never refuses to generate a biography of the leaders in our list.
-Comparison of GPT-3.5 & GPT-4
-The qualitative difference between GPT-3.5 and GPT-4 is also striking.
-For example, while GPT-3.5 struggles to even generate fluent and coherent texts in Bengali, GPT-4 Bengali generations are mostly fluent and accurate, however the generations were short and lacking specific details.
-One interesting example regarding difference between GPT-3.5 and GPT-4 comes from Arabic.
-In GPT-3.5, when asked about Pakistani former president Mamnoon Hussain, GPT-3.5 hallucinates a biography non-existent "ambitious young man from Saudi Arabia", while GPT-4 rightfully identifies and writes the biography of former Pakistani president.
-GPT-4 also refused to answer more often than GPT-3.5; For example, in Chinese, it refuses to answer the biography of Sushir Koirala and claims there is no famous person of this name, however goes on to generate a fictional biography of an Indian person (with a disclaimer).
-GPT-4, too, does not always refrain from answering, when it doesn’t know.
-For example, when asked about Ollanta Humala in Chinese, it goes on to describe Ollantay Tambo, an archeological site in Peru.
-5 Discussion & Limitations
-In this study, we introduce a unique approach to examining the distribution of geo-culturally based knowledge using atomic factuality in long-form LLM generation.
-We observe that LLMs exhibit a Western-centric bias, which remains evident even in languages from non-Western regions.
-This discovery prompts intriguing inquiries regarding the representation of multilingual knowledge within LLMs.
-For instance, do LLMs learn to represent more facts about South America in Bengali by being exposed solely to content in English or Spanish?
-Or, to what extent is multilingual data required to acquire factual knowledge in a specific language?
-However, due to the proprietary nature of models like GPT-4, answering these questions proves to be challenging.
-Future research should delve deeper into this issue with language models that also have open datasets.
-Recent studies such as Dolma Soldaini et al.
-(2024) and OLMo Groeneveld et al.
-(2024) represent a step forward in this regard, although their focus is limited to monolingual English corpora and models.
-One limitation of our current study pertains to small sample bias.
-We examine only a single sample from each country, which may not provide a comprehensive representation, especially considering that automated FActScore sometimes yields slight overestimations or underestimations.
-Consequently, caution is needed when interpreting individual scores.
-For instance, assertions such as "In Korean, the biography of the prime minister of Malaysia is 5% more accurate than that of the Romanian president" should be made with caution.
-In this paper, we present all findings as averages across languages, continents, or regions to mitigate sample bias.
-Future research should include multiple individuals, not only political leaders, from each country to alleviate this limitation.
-Another limitation of our methodology is the varying durations national leaders have been in power, potentially biasing internet corpora in their favor.
-One solution could be to prompt models to generate a “country biography” or “capital biography” (e.g., “Write me an essay on Addis Ababa”).
-However, we opted not to pursue this in our study, as the original FActScore research did not include a human evaluation baseline for non-biographical domains.
-6 Related Work
-LLM Factuality Evaluation
-The assessment of factual accuracy in natural language texts predates the advent of LLMs Guo et al.
-(2022).
-Our Multi-FAct pipeline follows the previous fact evaluation research, which involves validating claims based on external resource such as Wikipedia article Thorne et al.
-(2018); Krishna et al.
-(2022); Zhong et al.
-(2019) or Google search Chern et al.
-(2023).
-Our work is heavily inspired and based on Min et al.
-(2023), which suggests to measure the factuality of a long-form generated output of LLMs by breaking it down into atomic facts.
-Another line of work focuses on evaluating factual accuracy solely based on model output or internal states, eliminating the need for external knowledge bases like Wikipedia Manakul et al.
-(2023); Azaria and Mitchell (2023); Dhuliawala et al.
-(2023).
-While such methods offer resource efficiency, they often sacrifice accuracy compared to approaches leveraging external resources for fact verification.
-Multilingual Factuality Evaluation
-In the realm of multilingual factuality, Gupta and Srikumar (2021) presents the X-Fact dataset for multilingual fact-checking.
-Similarly, Thorne et al.
-(2018) have developed frameworks for multilingual fact-extraction and verification evaluation.
-Additionally, Aharoni et al.
-(2022) introduce mFACE, a dataset aimed at evaluating multilingual factual consistency.
-While these contributions significantly advance the understanding of multilingual factuality, they primarily focus on the assessment of factual accuracy in existing articles or benchmarks rather than examining the factual accuracy in the generative outputs of models.
-In this vein, Kang et al.
-(2024) explore the efficacy of existing metrics designed to detect hallucinations of LLM-generations, proposed for English texts, in a multilingual setting and point out they fall short in multilingual contexts.
-Our study extends the endeavor of developing a factuality evaluation metric that can be applied in multiple languages by introducing a Multi-Fact pipeline.
-Geo-culture biases of LLMs
-There has been considerable research regarding cultural and linguistic bias in LLMs Hovy and Yang (2021); Cao et al.
-(2023); Hershcovich et al.
-(2022); Huang and Yang (2023); Jin et al.
-(2023).
-A concurrent work Mirza et al.
-(2024) examines geographic knowledge gaps in LLMs and finds that GPT models show higher factuality in the Global North over the Global South, consistent with our findings.
-Manvi et al.
-(2024) also reach a similar conclusion.
-However, none of the studies concerns free-form generation with non-English languages.
-7 Conclusion & Future Work
-In this paper, we introduce a simple pipeline to evaluate factuality of LLMs in multilingual settings and demonstrate its effectiveness with the national leaders’ biographies in 80 countries around the world.
-Our findings reveal that multilingual models, such as GPT-4, tend to produce more factual content in high-resource languages, with notably longer outputs in English.
-Additionally, we observe a pervasive Western-centric bias across all languages; on average, even Asian and African languages demonstrate better performance for American and European leaders.
-However, a more detailed geographical analysis indicates that languages still reflect some aspects of their local geographic distribution, albeit to a lesser degree.
-While factuality might be an interesting way to look into a model’s knowledge about a certain person, our current implementation treats every fact equally.
-FActScore does not differentiate between the informational value of facts.
-For instance, the statements “Person X attended MIT” and “Person X attended a famous university” are treated as equivalent, though the former is obviously more informative and useful.
-Future research should aim to distinguish between specific, valuable facts and generic, less informative ones.
-Another area for future investigation could involve examining the consistency of model-generated facts across different languages; for example, whether the model’s depiction of Barack Obama’s childhood in English aligns with its portrayal in Chinese, or whether it keeps its focus solely on his political endeavors.
-Appendix A Appendix
-A.1 List of Countries and Presidents
-Table 2 lists all the names of the countries, presidents and their Geographic locations. Please note that, we changed Australia and New Zealand to Oceania in the UN geoscheme for plotting purpose.
+Appendix E List of Countries and Presidents
+Table 8 lists all the names of the countries, presidents and their Geographic locations. Please note that, we changed Australia and New Zealand to Oceania in the UN geoscheme for plotting purposes.
 Country Continent Region President
@@
 Switzerland Europe Western Europe Simonetta Sommaruga
-Table 2:
-A.2 Multi-FAct Implementation Details
-We replicated the original FActScore implementation and replaced InstructGPT for atomic fact generation with Mistral-7B.
-We also used Mistral 7B for the retrieval-based verification step.
-The final version uses retrieval + Mistral-7B + NPM.
-NPM threshold was set to 0.3.
-A.2.1 Replication of original FActScore
-We provide a comparison between our implementation with the results reported in the paper.
-For all the following experiments, we used a subset of ChatGPT-generation human-annotated factual data from the original paper.
-Metric Human FS FS (ours) Error (ours) Error (Original)
-Value 0.626 $\pm$ 0.238 0.637 $\pm$ 0.207 -0.011 $\pm$ 0.129 -0.0383 $\pm$ 0.121
-Table 3: Summary of Replication.
-Error refers to the difference between human FActScore and FActScore determined by our method.
-Error (Original) refers to the error by the best-performing method reported on the paper.
-We also report the standard deviation alongside each metric.
-Note that the Human FActScore value in the table 3 is different from what is provided in the paper as we averaged over our subset only.
-We replicated the best-performing method (retrieval + NPM) on this dataset reported in the paper and marked it as Error (Original).
-Our higher accuracy is likely due to Mistral 7B’s better capabilities and our better hyperparameter tuning.
-Please note that there is some stochasticity in the automated evaluation of FActScore , which is why we need to be cautious of individual scores, and only work with scores averaged over a large number of samples.
-A.2.2 Evaluation of Multi-FAct Pipeline
-The basis of our pipeline relies crucially on the GPT3.5-based translation (Step 3 in figure 2).
-It is imperative to see whether this translation step hurts the automated estimation of FActScore in a multilingual setting.
-Evaluation Method:
-The basis of our evaluation method is that the factuality of a text is translation invariant, ie, factual knowledge in a text should not change when it is translated to other languages.
-To evaluate this, we first take the human-annotated examples provided in original FActScore , and translate them into four languages using GPT4.
-We use GPT4 as previous studies Jiao et al.
-(2023) revealed GPT4 to be a good translator and better than GPT3.5 and commercially available machine translators.
-Manual inspection by native speakers of a subset of our languages also confirmed this.
-Language/Dataset FActScore
-Human eval (EN) 0.626 $\pm$ 0.238
-EN 0.637 $\pm$ 0.207
-KOR 0.625 $\pm$ 0.211
-BN 0.600 $\pm$ 0.216
-FR 0.638 $\pm$ 0.200
-SW 0.654 $\pm$ 0.196
-Table 4: FActScore evaluation Across Languages.
-EN refers to FActScore evaluation on original English human annotation; other languages refer to FActScore evaluation on GPT4 translation of the human annotation.
-Figure 9: Correlation of FActScore evaluated on translated human-annotated data.
-We then applied our multilingual FActScore pipeline (Figure 2) on this GPT4 translated text.
-The premise is that if our pipeline (which involves GPT3.5 translation [^6] We did not use GPT4 translation in Multi-FAct pipeline to reduce API costs) provides similar FActScore as the original English annotated data, it might provide good evidence that having a translation step in our pipeline doesn’t hurt FActScore estimation significantly.
-The results are provided in Table 4. We see that running Multi-FAct on non-English data does not significantly degrade FActScore estimation.
-We also note on the bottom row of Figure (9) that the correlation between FActScore evaluated in English and FActScore evaluated in other languages for the texts are very high; and all of them are very closely correlated with human-judgement.
-A.3 Additional Results
-Regional Breakdown (GPT4):
-Figure 10 shows the regional breakdown for the other languages that were not included in Figure 5.
-(a) Spanish
-(b) German
-(c) Swahili
-(d) French
-(e) Bengali
-Figure 10: Fine-grained Geographic Distribution for Remaining Languages
+Table 8: List of Countries and Presidents Used in Our Experiment